# On the Resistance Conjecture

### v1

@@ Abstract.
We thank Mathav Murugan and Riku Anttila for extensive discussions on the problem and for asking great questions, which lead to identifying the Whitney blending technique. Both of them had observed earlier the importance of writing the cutoff Sobolev as a Poincaré inequality and that the cutoff Sobolev inequality would follow from the log-Caccioppoli inequality for functions $f$ with $f=0$ outside $2B$. Anttila was the first to publish an approach to cutoff Sobolev based on these ideas and proved partial results that came quite close [5]. After the first version of this draft, Murugan also shared us an unpublished draft which he had written and shared with colleagues in 2018, where he proved a version of cutoff Sobolev for $f$ vanishing outside $2B$. Murugan also had the early idea of decomposing a function $f$ into two parts. The missing piece that we add to the resolution here is the Whitney blending metbod, along with a few smaller but crucial technical observations. We thank Pekka Koskela for many lectures years ago teaching us a very similar method, and Juha Kinnunen for advocating for developing a theory applicable to all exponents $p$. We also thank Peking University for hosting the author in Fall 2025, and for the opportunity to give a sequence of lectures there, during which some of the ideas of the present paper matured. The author was partially supported by the Research Council of Finland via the project GeoQuantAM: Geometric and Quantitative Analysis on Metric spaces, grant no. 354241.
@@ 1.1. Main results
The resistance conjecture was posed by Grigor’yan, Hu and Lau in [32, Conjecture 4.16], [33, p. 1494], [5, Conjecture 4.28]. It asks if volume doubling, resistance bounds and a Poincaré inequality suffice to prove a parabolic Harnack inequality for general metric measure Dirichlet spaces. Similar, and less specific, general questions seeking to find minimal conditions that characterize parabolic Harnack inequalities were posed in [60, Open problem II] and [8, Remark 3.17]. In [11, 10], it was shown that Poincaré inequalities, doubling and the cutoff Sobolev inequality were equivalent to a parabolic Harnack inequality. The cutoff Sobolev inequality was quite mysterious and numerous authors commented on the difficulty of verifying it for explicit spaces. Further, it was unclear what its geometric content was. Here, we remove the cutoff Sobolev inequality from the characterization by showing that it is implied by volume doubling, Poincaré and upper capacity bounds thus leading to a full generalization of the seminal result of Grigor’yan and Saloff-Coste in [30, 75].
@@
Let $(X,d,\mu,\operatorname{\mathcal{E}},\operatorname{\mathcal{F}})$ be a geodesic metric measure Dirichlet space, then the following are equivalent:
@@
    $(X,d,\mu,\operatorname{\mathcal{E}},\operatorname{\mathcal{F}})$ satisfies volume doubling of $\mu$, Poincaré inequality $\operatorname{PI}(\beta)$ and capacity upper bound $\operatorname{\rm Cap}_{2}(\beta)$
For definitions of parabolic Harnack inequalities, two-sided capacity bounds ([50, Definition 4.2]), Heat kernel estimates and metric measure Dirichlet spaces, see e.g. [32, 50, 28] - we will not use these directly in our proofs and thus do not discuss them in more detail here; see [36, 60, 32, 80] for clear surveys and [50, 5] for well-written overviews on the literature and background. This result is a combination of a fairly large array of papers, and we recommend to see [50, Theorem 4.5] for a more detailed overview of the proof of the known parts. We briefly recall the history. In $\mathbb{R}^{n}$, elliptic and parabolic Harnack inequalities were established by Moser [66, 67]. This was extended to manifolds with non-negative Ricci curvature in [24, 62, 83]. If $\beta=2$, then $\operatorname{\rm Cap}_{2}(\beta)$ is automatically satisfied and the result is known from earlier works, in the manifold case [30, 75, 34]. The present version, with the addition of the cutoff Sobolev inequality was largely given in [10], while the case for graphs was studied in [11, 35].
@@
If a metric measure Dirichlet space satisfies volume doubling, $\operatorname{\rm Cap}_{p}(\beta)$ and a Poincaré inequality, then its martingale dimension is finite and it admits a measurable Riemannian structure in the sense of [46, Theorems 3.1 and 3.4].
@@
The main technique and new idea in the proof of Theorem 1.4 is a blending technique: if $f,g\in\operatorname{\mathcal{F}}_{p}$ and $B$ is a ball, then then there exists a function $h\in\operatorname{\mathcal{F}}_{p}$ s.t. $h|_{B}=f|_{B}$ and $h|_{X\setminus 2B}=f|_{X\setminus 2B}$. Such a blending is known to be possible assuming the cutoff Sobolev inequality, but the novelty here is that one can prove it directly from capacity bounds and a Poincaré inequality by adapting an old method of Whitney [81] via extension domains in [48], [37]. Such Whitney covers have been frequently used (in particular [70]), but not in the study of the cutoff Sobolev inequality. This method is general, and appears useful in other problems as well. Indeed, Mathav Murugan has already adapted this method to resolve a non-local version of the resistance conjecture in [72].
@@
We end this overview with an open problem. The Poincaré inequality is in some cases almost as difficult to prove as the cutoff Sobolev inequality. However, in many settings one can show the following balled-capacity estimate.
@@
We consider Theorem 1.4 as partial evidence towards a positive answer. Entirely new and different methods seem to be needed for its proof, however. If true, this would lead to an even simpler condition for elliptic and parabolic Harnack inequalities, and answer another problem [12, Problem 7.7]. Currently, a method of Bonk and Kleiner [15, Proposition 3.1] and Heinonen and Koskela [41, Corollary 5.13] as adapted by Murugan and Shimizu [68, Sections 3 and 4] indicates that this is true under the additional constraint of $\beta>Q-1$, when $\mu(B(x,r))\sim r^{Q}$. The general case is now known. It would help in proving Harnack inequalities for a variety of symmetric constructions for which it is presently not known, see e.g. [4, 20, 25, 26] for some examples of such spaces. We next discuss the literature on the cutoff Sobolev inequality in more detail and give an outline the main proof ideas.
@@ 1.3. Key proof ideas
The second observation is a way to reduce effectively to the case similar to $f=0$ outside $2B$. Indeed, in an attempt to do so, one is led to show the following: there exists a function $h\in\operatorname{\mathcal{F}}_{p}$ s.t. $h|_{3/2B}=f|_{3/2B}$ and $h|_{X\setminus 2B}=f_{B}$. Such a function $h$ would at least capture a part of the function $f$ and have the desired boundary data. In this case, we could compute:
@@ 4.2. Discrete convolutions
Assume $\operatorname{\rm Cap}_{p}(\Psi)$. Let $h\in L^{p}(X)$ s.t. there is a constant $\Lambda$ and a finite measure $\nu$ and a $\delta>0$ s.t. for all balls $B\subset X$ with $\operatorname{rad}(B)\leq\delta$ we have
@@
We will show that $h_{k}\in\operatorname{\mathcal{F}}_{p}$ with the desired energy bound.
First, Let $\operatorname{\mathcal{B}}_{n,k}\subset\operatorname{\mathcal{B}}_{k}$, for $n\in\mathbb{N}$ be a sequence of increasing finite sets s.t. $\bigcup_{n\in\mathbb{N}}\operatorname{\mathcal{B}}_{n,k}=\operatorname{\mathcal{B}}_{k}$, and let
$h_{n,k}=\sum_{B\in\operatorname{\mathcal{B}}_{n,k}}h_{B}\phi_{B}.$
As a finite sum of functions in $\operatorname{\mathcal{F}}_{p}$, we have $h_{n,k}\in\operatorname{\mathcal{F}}_{p}$ for all $n,k\in\mathbb{N}$. Next, $h_{n,k}\leq H$ with $H=\sum_{B\in\operatorname{\mathcal{B}}_{k}}|h_{B}|\phi_{B}$. Moreover,
$\int|H|^{p}d\mu$ $\lesssim\int\sum_{B\in\operatorname{\mathcal{B}}_{k}}|h_{B}|^{p}|\phi_{B}|d\mu$ $\operatorname{\mathcal{B}}_{k}$ is locally finite
$\leq\sum_{B\in\operatorname{\mathcal{B}}_{k}}|h_{B}|^{p}\mu(2B)$ $\phi_{B}:X\to[0,1]$ has support in $2B$
$\lesssim\sum_{B\in\operatorname{\mathcal{B}}_{k}}|h_{B}|^{p}\mu(B)$ doubling
$\leq\sum_{B\in\operatorname{\mathcal{B}}_{k}}\int_{B}|h|^{p}d\mu$ Jensen
$\lesssim\int_{X}|h|^{p}d\mu$ $\text{$\operatorname{\mathcal{B}}_{k}$ is locally finite}.$
Thus, $H\in L^{p}(X)$ and since $h_{n,k}\to h_{k}$ pointwise, by dominated convergence $h_{n,k}\to h_{k}$ in $L^{p}(X)$.
Further,
@@
Finally, we estimate the energies of $h_{n,k}$. Let $A\subset X$ be a closed set and let $A_{k}=\{y\in X:d(y,A)\leq 2^{2-k}\Lambda\}$. First, if $B\cap B^{\prime}\neq\emptyset$ for $B,B^{\prime}\in\operatorname{\mathcal{B}}_{k}$, then $B^{\prime}\subset 4B$ and the doubling property and the Poincaré inequality imply that
@@
By applying these, we give a uniform bound for the energy of $h_{n,k}$:
$\Gamma_{p}\langle h_{n,k}\rangle(A)$ $\leq\sum_{B\in\operatorname{\mathcal{B}}_{k}}\Gamma_{p}\langle h_{n,k}\rangle(B\cap A)$ balls cover the set
$\leq\sum_{B\in\operatorname{\mathcal{B}}_{k}}\Gamma_{p}\langle\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{n,k}\end{subarray}}h_{B^{\prime}}\psi_{B^{\prime}}\rangle(B\cap A)$ locality
$\leq\sum_{B\in\operatorname{\mathcal{B}}_{k}}\Gamma_{p}\langle\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{n,k}\end{subarray}}(h_{B^{\prime}}-h_{B})\psi_{B^{\prime}}\rangle(B\cap A)$ partition of unity
$\lesssim\sum_{\begin{subarray}{c}B\in\operatorname{\mathcal{B}}_{k}\\ B\cap A\neq\emptyset\end{subarray}}\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{n,k}\end{subarray}}\Gamma_{p}\langle(h_{B^{\prime}}-h_{B})\psi_{B^{\prime}}\rangle(B)$ locally finiteness
$\lesssim\sum_{\begin{subarray}{c}B\in\operatorname{\mathcal{B}}_{k}\\ B\cap A\neq\emptyset\end{subarray}}\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{n,k}\end{subarray}}|h_{B^{\prime}}-h_{B}|^{p}\frac{\mu(B^{\prime})}{\Psi(B^{\prime})}$
@@
$\lesssim\sum_{\begin{subarray}{c}B\in\operatorname{\mathcal{B}}_{n,k}\\ B\cap A\neq\emptyset\end{subarray}}\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{n,k}\end{subarray}}|h_{B^{\prime}}-h_{B}|^{p}\frac{\mu(B(x,r))}{\Psi(x,r)}$ doubling
@@
By using $A=X$, we get $\Gamma_{p}\langle h_{n,k}\rangle(X)\lesssim\nu(X)$. Sending $n\to\infty$ and then $k\to\infty$ and using lower semi-continuity, we get $h\in\operatorname{\mathcal{F}}_{p}$ with $\Gamma_{p}\langle h\rangle(X)\lesssim\nu(X)$. For general closed sets $A$, we get $\Gamma_{p}\langle h\rangle(A)\lesssim\nu(A_{k})$ for every $k\in\mathbb{N}$, and then sending $k\to\infty$ yields $\Gamma_{p}\langle h\rangle(A)\lesssim\nu(A)$. By Borel regularity, we get the claim.
@@ 4.4. Blending procedure
Subcase C.1) $\operatorname{rad}(B_{*})\leq\operatorname{rad}(B)$: In this case
@@
Subcase C.2) $\operatorname{rad}(B_{*})>\operatorname{rad}(B)$: This is the final case. Lemma 4.8 implies that for all $B^{\prime}\in\operatorname{\mathcal{B}}$ s.t. $B^{\prime}\cap B\neq\emptyset$

### v2

@@ Abstract.
We thank Mathav Murugan and Riku Anttila for extensive discussions on the problem and for asking great questions, which lead to identifying the Whitney blending technique. Both of them had observed earlier the importance of writing the cutoff Sobolev as a Poincaré inequality and that the cutoff Sobolev inequality would follow from the log-Caccioppoli inequality for functions $f$ with $f=0$ outside $2B$. Anttila was the first to publish an approach to cutoff Sobolev based on these ideas and proved partial results that came quite close [5]. After the first version of this draft, Murugan also shared us an unpublished draft which he had written and shared with colleagues in 2018, where he proved a version of cutoff Sobolev for $f$ vanishing outside $2B$. Murugan also had the early idea of decomposing a function $f$ into two parts. The missing piece that we add to the resolution here is the Whitney blending method, along with a few smaller but crucial technical observations. We thank Pekka Koskela for many lectures years ago teaching us a very similar method, and Juha Kinnunen for advocating for developing a theory applicable to all exponents $p$. We also thank Ryosuke Shimizu for a careful reading of the draft of the paper and for his comments that helped correct several inaccuracies. Peking University hosted the author in Fall 2025, and we thank it for the opportunity to give a sequence of lectures there, during which some of the ideas of the present paper matured. The author was partially supported by the Research Council of Finland via the project GeoQuantAM: Geometric and Quantitative Analysis on Metric spaces, grant no. 354241.
@@ 1.1. Main results
The resistance conjecture was posed by Grigor’yan, Hu and Lau in [32, Conjecture 4.16], [33, p. 1494], [5, Conjecture 4.28]. It asks if volume doubling, resistance bounds and a Poincaré inequality suffice to prove a parabolic Harnack inequality for general metric measure Dirichlet spaces. Similar, and less specific, general questions seeking to find minimal conditions that characterize parabolic Harnack inequalities were posed in [60, Open problem II] and [8, Remark 3.17]. In [11, 10], it was shown that Poincaré inequalities, doubling and the cutoff Sobolev inequality were equivalent to a parabolic Harnack inequality. The cutoff Sobolev inequality was quite mysterious and numerous authors commented on the difficulty of verifying it for explicit spaces. Further, it was unclear what its geometric content was. Here, we remove the cutoff Sobolev inequality from the characterization by showing that it is implied by volume doubling, Poincaré and upper capacity bounds thus leading to a conceptually simple generalizations of the seminal result of Grigor’yan and Saloff-Coste in [30, 75], simplifying the characterization in [10].
@@
Let $(X,d,\mu,\operatorname{\mathcal{E}},\operatorname{\mathcal{F}})$ be a metric measure Dirichlet space, then the following are equivalent:
@@
    $(X,d,\mu,\operatorname{\mathcal{E}},\operatorname{\mathcal{F}})$ satisfies volume doubling of $\mu$, Poincaré inequality $\operatorname{PI}(\beta)$ and capacity upper bound $\operatorname{\rm Cap}_{2}(\beta)$.
Notice that here, and below, we avoid any assumptions on the metric being geodesic. For definitions of parabolic Harnack inequalities, two-sided capacity bounds ([50, Definition 4.2]), Heat kernel estimates and metric measure Dirichlet spaces, see e.g. [32, 50, 28] - we will not use these directly in our proofs and thus do not discuss them in more detail here; see [36, 60, 32, 80] for clear surveys and [50, 5] for well-written overviews on the literature and background. This result is a combination of a fairly large array of papers, and we recommend to see [50, Theorem 4.5] for a more detailed overview of the proof of the known parts. We briefly recall the history. In $\mathbb{R}^{n}$, elliptic and parabolic Harnack inequalities were established by Moser [66, 67]. This was extended to manifolds with non-negative Ricci curvature in [24, 62, 83]. If $\beta=2$, then $\operatorname{\rm Cap}_{2}(\beta)$ is automatically satisfied and the result is known from earlier works, in the manifold case [30, 75, 34]. The present version, with the addition of the cutoff Sobolev inequality was largely given in [10], while the case for graphs was studied in [11, 35].
@@
If a metric measure Dirichlet space satisfies volume doubling, $\operatorname{\rm Cap}_{p}(\beta)$ and a Poincaré inequality $\operatorname{PI}_{p}(\beta)$, then its martingale dimension is finite and it admits a measurable Riemannian structure in the sense of [46, Theorems 3.1 and 3.4].
@@
The main technique and new idea in the proof of Theorem 1.4 is a blending technique: if $f,g\in\operatorname{\mathcal{F}}_{p}$ and $B$ is a ball, then there exists a function $h\in\operatorname{\mathcal{F}}_{p}$ s.t. $h|_{B}=f|_{B}$ and $h|_{X\setminus 2B}=f|_{X\setminus 2B}$. Such a blending is known to be possible assuming the cutoff Sobolev inequality, but the novelty here is that one can prove it directly from capacity bounds and a Poincaré inequality by adapting an old method of Whitney [81] via extension domains in [48], [37]. Such Whitney covers have been frequently used (in particular [70]), but not in the study of the cutoff Sobolev inequality. This method is general, and appears useful in other problems as well. Indeed, Mathav Murugan has already adapted this method to resolve a non-local version of the resistance conjecture in [72].
@@
We end this overview with an open problem. The Poincaré inequality is in some cases almost as difficult to prove as the cutoff Sobolev inequality. However, in many settings one can show quite easily the following balled-capacity estimate.
@@
In fact, for $p\neq 2$ it is unkown if this holds even for quite simple fractals, such as the generalized Sierpiński carpets. We consider Theorem 1.4 as partial evidence towards a positive answer. Entirely new and different methods seem to be needed for its proof, however. If true, this would lead to an even simpler condition for elliptic and parabolic Harnack inequalities, and answer another problem [12, Problem 7.7]. Currently, a method of Bonk and Kleiner [15, Proposition 3.1] and Heinonen and Koskela [41, Corollary 5.13] as adapted by Murugan and Shimizu [68, Sections 3 and 4] indicates that this is true under the additional constraint of $\beta>Q-1$, when $\mu(B(x,r))\sim r^{Q}$. The general case is not known. It would help in proving Harnack inequalities for a variety of symmetric constructions for which it is presently not known, see e.g. [4, 20, 25, 26] for some examples of such spaces. We next discuss the literature on the cutoff Sobolev inequality in more detail and give an outline the main proof ideas.
@@ 1.3. Key proof ideas
The second observation is a way to reduce effectively to the case similar to $f=0$ outside $2B$. Indeed, in an attempt to do so, one is led to show the following: there exists a function $h\in\operatorname{\mathcal{F}}_{p}$ s.t. $h|_{3/2B}=f|_{3/2B}$ and $h|_{X\setminus 2B}=f_{B}$. (Technically, when $f_{B}\neq 0$, and the space is unbounded, we would need to consider a local Sobolev space here.) Such a function $h$ would at least capture a part of the function $f$ and have the desired boundary data. In this case, we could compute:
@@ 4.2. Discrete convolutions
Assume $\operatorname{\rm Cap}_{p}(\Psi)$. Let $h\in L^{p}(X)$ have compact support s.t. there is a constant $\Lambda$ and a finite measure $\nu$ and a $\delta>0$ s.t. for all balls $B\subset X$ with $\operatorname{rad}(B)\leq\delta$ we have
@@
We will show that $h_{k}\in\operatorname{\mathcal{F}}_{p}$ with the desired energy bound. Since the support of $h$ is compact, this sum always has finitely many non-zero terms. Similarly, all the sums below will be finite in this same sense.
As a finite sum of functions in $\operatorname{\mathcal{F}}_{p}$, we have $h_{k}\in\operatorname{\mathcal{F}}_{p}$ for all $k\in\mathbb{N}$. Further,
@@
Finally, we estimate the energies of $h_{k}$. Let $A\subset X$ be a closed set and let $A_{k}=\{y\in X:d(y,A)\leq 2^{2-k}\Lambda\}$. First, if $B\cap B^{\prime}\neq\emptyset$ for $B,B^{\prime}\in\operatorname{\mathcal{B}}_{k}$, then $B^{\prime}\subset 4B$ and the doubling property and the Poincaré inequality imply that
@@
By applying these, we give a uniform bound for the energy of $h_{k}$:
$\Gamma_{p}\langle h_{k}\rangle(A)$ $\leq\sum_{B\in\operatorname{\mathcal{B}}_{k}}\Gamma_{p}\langle h_{k}\rangle(B\cap A)$ balls cover the set
$\leq\sum_{B\in\operatorname{\mathcal{B}}_{k}}\Gamma_{p}\langle\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{k}\end{subarray}}h_{B^{\prime}}\psi_{B^{\prime}}\rangle(B\cap A)$ locality
$\leq\sum_{B\in\operatorname{\mathcal{B}}_{k}}\Gamma_{p}\langle\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{k}\end{subarray}}(h_{B^{\prime}}-h_{B})\psi_{B^{\prime}}\rangle(B\cap A)$ partition of unity
$\lesssim\sum_{\begin{subarray}{c}B\in\operatorname{\mathcal{B}}_{k}\\ B\cap A\neq\emptyset\end{subarray}}\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{k}\end{subarray}}\Gamma_{p}\langle(h_{B^{\prime}}-h_{B})\psi_{B^{\prime}}\rangle(B)$ locally finiteness
$\lesssim\sum_{\begin{subarray}{c}B\in\operatorname{\mathcal{B}}_{k}\\ B\cap A\neq\emptyset\end{subarray}}\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{k}\end{subarray}}|h_{B^{\prime}}-h_{B}|^{p}\frac{\mu(B^{\prime})}{\Psi(B^{\prime})}$
@@
$\lesssim\sum_{\begin{subarray}{c}B\in\operatorname{\mathcal{B}}_{n,k}\\ B\cap A\neq\emptyset\end{subarray}}\sum_{\begin{subarray}{c}B^{\prime}\cap B\neq\emptyset\\ B^{\prime}\in\operatorname{\mathcal{B}}_{k}\end{subarray}}|h_{B^{\prime}}-h_{B}|^{p}\frac{\mu(B(x,r))}{\Psi(x,r)}$ doubling
@@
By using $A=X$, we get $\Gamma_{p}\langle h_{k}\rangle(X)\lesssim\nu(X)$. Sending $k\to\infty$ and using lower semi-continuity, we get $h\in\operatorname{\mathcal{F}}_{p}$ with $\Gamma_{p}\langle h\rangle(X)\lesssim\nu(X)$. For general closed sets $A$, we get $\Gamma_{p}\langle h\rangle(A)\lesssim\nu(A_{k})$ for every $k\in\mathbb{N}$, and then sending $k\to\infty$ yields $\Gamma_{p}\langle h\rangle(A)\lesssim\nu(A)$. By Borel regularity, we get the claim.
@@ 4.4. Blending procedure
Subcase B.1) $\operatorname{rad}(B_{*})\leq\operatorname{rad}(B)$: In this case
@@
Subcase B.2) $\operatorname{rad}(B_{*})>\operatorname{rad}(B)$: This is the final case. Lemma 4.8 implies that for all $B^{\prime}\in\operatorname{\mathcal{B}}$ s.t. $B^{\prime}\cap B\neq\emptyset$